# UDEEP: Edge-based Computer Vision for In-Situ Underwater Crayfish and Plastic Detection

### v1

@@ Abstract
Invasive signal crayfish have a detrimental impact on ecosystems. They spread the fungal-type crayfish plague disease (Aphanomyces astaci) that is lethal to the native white clawed crayfish, the only native crayfish species in Britain. Invasive signal crayfish extensively burrow, causing habitat destruction, erosion of river banks and adverse changes in water quality, while also competing with native species for resources and leading to declines in native populations. Moreover, pollution exacerbates the vulnerability of White-clawed crayfish, with their populations declining by over 90% in certain English counties, making them highly susceptible to extinction. To safeguard aquatic ecosystems, it is imperative to address the challenges posed by invasive species and discarded plastics in the United Kingdom’s river ecosystem’s. This article introduces the UDEEP, Underwater Deep learning EdgE computing Platform, for the detection of crayfish and plastic. The paper also presents two publicly available underwater datasets, annotated with sequences of crayfish and underwater plastic debris, captured in freshwater rivers under varying visibility conditions. The crayfish and plastic object detection training achieved the best accuracy of 0.9 mean Average Precision (mAP) @0.5 Intersection over Union (IoU) on You Only Look Once (YOLO) v5s. YOLO v8n achieved the highest precision at 92%. Subsequently, YOLO v5n achieved the best recall score overall at 93%. As for the amount current required for inference, YOLO v8s required the least amount of current for inference, 2548mA and YOLO v5n consumed the highest current at 3277mA.
The UDEEP platform can play a crucial role in environmental monitoring by performing on-the-fly classification of Signal crayfish and plastic debris while leveraging the efficacy of Artificial Intelligence (AI), Internet of Things (IOT) devices and the power of edge computing (i.e., NVIDIA Jetson Nano (NJN)). By providing accurate data on the presence, spread and abundance of these species, the UDEEP platform can contribute to monitoring efforts and aid in mitigating the spread of invasive species.
@@
Deep Learning (DL), Signal crayfish, Invasive Species, Plastics, image segmentation, YOLO v5, YOLO v8
@@
[NTU]organization=Nottingham Trent University,Department of Computer Science addressline=Clifton Campus, city=Nottingham, postcode=NG11 8NS, country=UK
[…]
[MAI]organization=Mediprospects AI, addressline=5-7 High Street, city=London, postcode=E13 0AD, country=UK
{graphicalabstract}
Research highlights {highlights}
UDEEP was developed for automated crayfish and plastic monitoring.
YOLO v5s,5n,8s and 8n models were trained and evaluated on an NJN board.
The models accurately detected crayfish and plastics in images under varying underwater conditions.
Two publicly available underwater datasets are introduced, annotated with sequences of Signal crayfish and plastic debris.
@@ 1 Introduction
Invasive species exert substantial impacts on ecosystems globally, comparable to the effects of pollutants such as discarded plastic. Extensive information exists about these species, including their characteristics, invaded regions, invasion pathways, and impacts Tekman et al. (2022); Hobbs (2000). However, there is still a need for comprehensive assessments that consider multiple aspects of the invasion process simultaneously Bernery et al. (2022); Woods et al. (2016).
@@
Invasive signal crayfish (Pacifastacus Leniusculus) pose a threat to native species and habitats, causing substantial economic consequences. These invasive crayfish out-compete and prey on native white-clawed crayfish (Austropotamobius Pallipes), causing population declines and even extinctions Mathers et al. (2016). The white-clawed crayfish is a keystone species found in freshwater bodies in Western Europe Mirimin et al. (2022). However, its populations have experienced significant declines due to the presence of a contagious fungal–type disease (Aphanomyces Astaci) originating from North America, competition with non-indigenous crayfish species, and habitat deterioration Mirimin et al. (2022). Additionally, their burrowing activities result in habitat degradation, destabilising riverbanks and disrupting aquatic vegetation. The economic impact of these damages includes the loss of biodiversity, reduced productivity in fisheries, and the need for habitat restoration, which can require substantial financial investments Faller et al. (2016). Access to accurate and comprehensive data on these invasions is crucial Bernery et al. (2022).
In the process of detecting and recognising aquatic animals, the initial step involves acquiring data and gathering information. However, due to the complexity and dynamicist of the aquatic ecosystem, collecting underwater images can be relatively difficult and even dangerous Li et al. (2023). Nevertheless, advancements in technology have made it possible to collect fine-scale data using unmanned systems and a variety of sensors, such as physical and environmental data, imagery, and bathymetry data. Bathymetry is the measurement of depth of water in rivers, lakes or oceans to produce depth maps. These can then be used to characterise details of animal behaviour and their interactions within the ecosystem Ling et al. (2016). DL is also being used to identify and locate aquatic species in raw images, videos, and audio recordings for bio-monitoring purposes Goodwin et al. (2022).
@@
Use of cameras for biodiversity monitoring has less impact on biodiversity and can cover larger areas and for longer periods compared to traditional observation methods, that involve observers walking in transects. Data that can be collected includes abundance, density, species richness and proportion of area occupied Stephenson (2020). Additionally, Light Detection and Ranging (LIDAR) detection can be added to measure ranges, temperature, light intensity sensors can be added to provide a more holistic dataset.
This paper introduces the Underwater UDEEP which combines various techniques, including background subtraction for selective cropping of regions of interest in high-resolution images for the detection of crayfish and underwater plastic debris. The platform plays a pivotal role in detecting and estimating the density of Signal crayfish and plastic debris by leveraging the efficacy of AI, IOT devices and the power of edge computing (i.e., NJN). By delivering precise data regarding the presence, distribution, and population density of these species and discarded plastics, the UDEEP platform significantly contributes to monitoring endeavours and facilitates efforts to mitigate the proliferation of invasive species and environmental pollution by plastic debris.
The remainder of this article includes the literature review in Section 2, the description of the computational experiments for training, testing and validation with YOLO variants 5s and 5n, 8s and 8n in Section 3, the results of the computational experiments Section 4, and finally the conclusions and future work in Section 5
@@ 2 Related Work
Recent years have seen a significant increase in the interest in processing underwater images Mittal et al. (2022). The interest is driven by the need to study and monitor aquatic plants and animals, which has applications in marine biology, economy, and biodiversity management. Understanding the behaviour and numbers of aquatic species is crucial for various purposes, including protecting endangered species, analysing differences in species, and early detection of climatic events like pollution and global warming. For instance, Plankton, which produces over 80% of the world’s oxygen Witman and Witman (2021), plays a critical role in the ocean’s food chain and atmosphere-ocean connection. Monitoring plankton levels is essential, as both low and excessive levels can have harmful consequences. Furthermore, organisms like Posidonia Oceanic, which thrive in clean water, are vital for biodiversity, beach erosion reduction, and water quality enhancement Mittal et al. (2022). Image processing complements other techniques like physiochemical water analysis and sonar-based detection in understanding the impact of global warming and human activity on marine life.
@@ 2.1 The challenges of underwater monitoring
Underwater imaging contends with substantial energy loss, resulting in diminished and inconsistent illumination and visibility, particularly in the deeper expanses Mittal et al. (2022). The situation is further exacerbated by the presence of freshwater and ocean currents, intensifying the degradation of visibility. Impurities, suspended solids, low contrast, and compromised edges and details further impact underwater images. The distortion of colour based on distance, caused by non-uniform spectral propagation, adds another layer of complexity. Additionally, underwater conditions present intricate and uncontrolled backgrounds, posing challenges in segregating subjects such as sea cucumbers and shellfish from their surroundings. The unbridled movement of fishes in three-dimensional space, often concealed behind objects, compounds the difficulty in accurately determining their pose, orientation, and size. The scarcity of large-scale, publicly available datasets for underwater images, coupled with imbalances, necessitates the incorporation of data augmentation techniques.
@@ 2.2 Early advances in underwater bio-diversity monitoring
In the early 90’s, Strachan et al. (1990) laid the groundwork for automated fish grading based on size and colour, utilising image analysis to estimate fish shape and colour. This innovative approach allowed for the differentiation of haddock fish stocks from the Rockall Plateau, located midway between Greenland and Iceland and the North Sea. The method involved digitised Red, Green, and Blue (RGB) images captured from beneath a conveyor belt, processed using bespoke software written in the C programming language.
The study marked a significant advancement in distinguishing between fish species. Building on the pioneering work, Storbeck and Daan (2001) conducted further research in 2001 by integrating computer vision with a neural network, expanding on the foundations set by Strachan et al. (1990).
Both studies required the fish to be laid flat on a conveyor belt moving at a specific speed, with a vertical arrangement of a laser source and camera above the belt. However, the classification method, while groundbreaking, proves impractical for underwater environments.
@@ 2.3 Modern application of computer vision for underwater bio-diversity monitoring
A more contemporary perspective is presented in the work of Zhang et al. Zhang et al. (2021), showcasing a modern application of computer vision for underwater object detection. Employing the YOLO v4 model, the authors achieved an mAP of 92.65% at a processing speed of approximately 45 Frames per Second (FPS). Despite the success, it’s crucial to note that the achievement relied on a high-end NVIDIA graphics card for accelerating the YOLO v4. However, deploying such hardware alongside sensors in physical underwater settings becomes unfeasible due to size and power constraints.
In response to the constraint, it is advisable to explore embedded systems as viable alternatives. Such platforms present a more pragmatic deployment option for sensors, ensuring the practical implementation of advanced computer vision techniques in underwater environments. As underwater computer vision emerges as a burgeoning field, studying analogous works becomes imperative to identify potential gaps and opportunities that can be leveraged to advance knowledge in this domain.
@@
For instance, Hegde et al. (2021) successfully implemented a trained MobileNetV2 model capable of running on a Raspberry Pi, achieving an impressive precision of 98% for mackerel. While acknowledging that a YOLO model could offer faster results, they deemed it unfeasible due to hardware limitations. Consequently, their image capture and processing occurred at a rate of once every 2 seconds, equivalent to about 0.5 FPS. This limitation could prove critical in scenarios involving Underwater Autonomous Vehicles, where objects of interest might traverse the sensors before processing completion.
@@
The incorporation of a YOLO v5 model on the NJN platform is set to enhance performance, delivering a finely tuned solution tailored for the target applications. Noteworthy is the study by Hegde et al. (2021), which had a limited implementation of comprehensive image pre-processing methods, mainly focusing on cropping. In contrast, our UDEEP algorithm surpasses these limitations by integrating background subtraction which is a technique enables to extract moving objects (foreground) from the background, as discussed by Huang et al. (2017) and Kalsotra and Arora (2022).
@@
The sophisticated enhancement is anticipated to streamline object identification, potentially fortifying our algorithm’s efficacy in detecting small objects Hegde et al. (2021). Importantly, our dataset boasts an extensive collection of [450 500] photos per each category, significantly contributing to the robustness of our approach. In a comparable study, Wang et al. (2020) employed YOLO nano on an NJN to accelerate the algorithm at the edge. The model used by Wang et al. Wang et al. (2020) demonstrated the capability to detect underwater objects at an impressive frame rate of 8 FPS, accompanied by a notable mAP of approximately 74.8%. Notably, a discernible trade-off emerges between runtime and accuracy when juxtaposed with the work of Hegde et al. (2021). The strategic use of image pre-processing might provide an avenue for simpler, faster models to attain higher accuracy in underwater object detection scenarios.
@@
Chen et al. (2021) presented a methodology similar to the one employed in this work, utilising data augmentation to enhance the capabilities of the object detection model. Wei et al. (2020) proposed mosaic augmentation as use to augment the dataset to enhance the detection of smaller objects in cluttered environments. Moreover, Wei et al. Wei et al. (2020) introduced an innovative adaptation, involving the calculation of gains on the RGB channels to modify and adjust images, accommodating variations in underwater light levels. The UDEEP distinguishes itself through the deliberate exclusion of the background. This intentional omission is crafted to emphasise the salient features within the image, establishing our methodology as distinct from approaches that encompass the entire scene.
@@
Considerable efforts have been directed towards advancing efficient autonomous underwater Underwater Autonomous Vehicles (UAVs) employing machine learning algorithms for debris detection. Leveraging the computational power of an NVIDIA Jetson TX2 Graphics Processing Unit (GPU), Fulton et al. (2019) conducted comprehensive experiments involving transfer learning on various models, including YOLO v2, Tiny-YOLO, Faster R-CNN, and Single Shot Detector (SSD). In their approach, the author fine-tuned each network individually for R-CNN, SSD, and Tiny-YOLO. However, for YOLO v2, an additional step involved training using transfer learning. The models were specifically tailored for underwater plastic debris detection, and the results indicated that Faster R-CNN achieved the highest accuracy, albeit with longer inference times. YOLO v2 struck a balance between speed and accuracy, SSD exhibited the best inference time, and Tiny-YOLO emerged as the top performer in terms of overall performance.
@@
Saleh et al. (2024) assert that automated fish phenotyping, encompassing parameters such as fish length, weight, size, and mortality, provides invaluable insights into understanding the comprehensive profile of classified fish. The task of individually counting fish is notably resource-intensive, demanding skilled labour for hours, thereby contributing to elevated project costs and extended delivery timelines. In a related domain, Ranjan et al. (2023) introduced a real-time mortality monitoring system based on YOLO v7, achieving a commendable mAP of 93.4% and an F1 score of 0.89. The UDEEP builds upon this foundation, not only incorporating these crucial findings, but also extending the phenotyping metrics. Specifically, we introduce age estimation, distinguishing between adult and juvenile crayfish.

### v2

@@ Abstract
Invasive signal crayfish have a detrimental impact on ecosystems. They spread the fungal-type crayfish plague disease (Aphanomyces astaci) that is lethal to the native white clawed crayfish, the only native crayfish species in Britain. Invasive signal crayfish extensively burrow, causing habitat destruction, erosion of river banks and adverse changes in water quality, while also competing with native species for resources leading to declines in native populations. Moreover, pollution exacerbates the vulnerability of White-clawed crayfish, with their populations declining by over 90%. To safeguard aquatic ecosystems, it is imperative to address the challenges posed by invasive species and pollution in aquatic ecosystem’s.
This article introduces the Cognitive Edge Device (CED) computing platform for the detection of crayfish and plastic. It also presents two publicly available underwater datasets, annotated with sequences of crayfish and aquatic plastic debris. Four You Only Look Once (YOLO) variants were trained and evaluated for crayfish and plastic object detection. YOLO v5s achieved the highest detection accuracy, with an mAP@0.5 of 0.90, and achieved the best precision at 0.93. YOLO v8s exhibited the best recall at 0.84 and the lowest validation box loss of 0.95. While YOLO v5n and YOLO v8n also performed well with mAP@0.5 scores of 0.87 and 0.86, respectively, the YOLO v5 variants generally showed higher residual errors and notably slower convergence, with YOLO v5n being the slowest to stabilise. For real-time embedded testing on the NVIDIA Jetson Orin Nano (NJON), YOLO v8n proved to be efficient on the Graphics Processing Unit (GPU). It recorded the fastest inference time at 0.29 seconds and the lowest energy consumption at 0.09Joules (J), YOLO v5n took 0.48 seconds and consumed 2.43J. Despite not leading in all individual detection metrics, YOLO v8n’s balanced mAP@0.5, fastest speed and minimal power draw makes it most suited for deployment on resource-constrained platforms like the NJON, where a compromise between accuracy and operational efficiency is required. Central Processing Unit (CPU) inference times were notably slower and energy consumption was higher across all models.
The CED platform can play a crucial role in environmental monitoring by performing on-the-fly detection of Signal crayfish and plastic debris while leveraging the efficacy of Artificial Intelligence (AI), Internet of Things (IOT) devices and the power of edge computing (i.e., NJON). By providing accurate data on the presence, spread and abundance of these species, the platform can contribute to monitoring efforts and aid in mitigating the spread of invasive species.
@@ 1 Introduction
Invasive species cause substantial damage to environmental ecosystems globally, this is comparable to the effects of pollutants such as discarded plastic. Extensive information exists about these species, including their characteristics, invaded regions, invasion pathways, and impacts [tekman2022impacts, hobbs2000invasive]. However, there is still a need for comprehensive assessments that consider multiple aspects of the invasion process simultaneously [bernery2022, woods2016towards].
@@
Invasive signal crayfish (Pacifastacus Leniusculus) pose a threat to native species and habitats, causing substantial economic consequences. They out-compete and prey on native white-clawed crayfish (Austropotamobius Pallipes), causing population declines and even extinctions [MATHERS2016207]. The white-clawed crayfish is a keystone species found in freshwater bodies in Western Europe [mirimin2022]. However, its populations have experienced significant declines due to the presence of a contagious fungal–type disease (Aphanomyces Astaci) originating from North America, competition with non-indigenous crayfish species, and habitat deterioration [mirimin2022]. Additionally, their burrowing activities result in habitat degradation, destabilising riverbanks and disrupting aquatic vegetation. The economic impact of these damages includes the loss of biodiversity, reduced productivity in fisheries, and the need for habitat restoration, which can require substantial financial investments [FALLER20161190]. Access to accurate and comprehensive data on these invasions is crucial for informed decision-making in environmental management. [bernery2022].
Machine Learning (ML) has revolutionised various fields, including environmental monitoring, by enabling real-time data analysis and decision-making. In aquatic environments, there is an increasing demand for automated systems capable of performing accurate and timely classification of species, pollutants, and underwater infrastructure conditions [Juan_and_Xu]. Traditional methods of aquatic monitoring, which often rely on manual data collection and expert analysis, are time-consuming, resource-intensive, and limited in their scalability. The integration of ML, IOT devices, and edge computing offers a promising solution to these challenges, allowing for faster, more efficient, and cost-effective monitoring systems [Goodwin_Morten].
@@
Use of cameras for biodiversity monitoring has less impact on biodiversity and can cover larger areas and for longer periods compared to traditional observation methods, that involve observers walking in transects in challenging environments. Data that can be collected includes abundance, density, species richness and proportion of area occupied [STEPHENSON202036]. Additionally, various sensors can be added to collect more holistic datasets. For example Light Detection and Ranging (LIDAR) sensors to measure distance, temperature sensors for temperature measurements and light intensity sensors.
[biber2013] concluded that conventional methods for monitoring invasive species and environmental pollution typically involve manual sampling, which is resource-intensive, slow and unable to provide real-time information required for timely intervention strategies. Thus, environmental conditions monitoring requires innovative approaches that can provide cost-effective, continuous, automated, and scalable monitoring solutions.
To address the challenges of monitoring invasive species and plastics pollution, this paper introduces CED, a novel integration of AI, IOT, and edge computing technologies. CED enhances real-time detection of invasive species and plastic debris, providing actionable insights to support conservation efforts and environmental management.
The key contributions in this article are below:
  1. 1.
    A significant bottleneck in developing effective underwater AI solutions is the lack of diverse and well-annotated datasets. Our motivation includes contributing to the research community by providing two publicly available underwater datasets, annotated with sequences of crayfish and aquatic plastic debris. These datasets, which feature challenging underwater conditions, aim to foster further research and development in this critical domain by providing standardised benchmarks for model training and evaluation.
  2. 2.
    Traditional monitoring methods, such as manual sampling and visual surveys, are labor-intensive and prone to observer bias. Traditional monitoring methods are also limited in spatial and temporal resolution, which makes it challenging to track species like the invasive signal crayfish or detect the distribution of plastic debris effectively. CED addresses these limitations by using state-of-the-art deep learning models, specifically YOLO-based object detection frameworks, to automate the identification of target species and debris in real time.
  3. 3.
    Monitoring efforts often rely on centralised systems that collect data remotely and process it later, introducing delays in decision-making. The latency can result in missed opportunities for timely intervention, such as rapid response to invasive species or plastic pollution events. CED integrates edge computing to process data locally, allowing instant analysis and on-the-fly decision-making.
  4. 4.
    Manual monitoring programs often incur high costs due to personnel requirements and extensive fieldwork. Additionally, traditional equipment for automated monitoring can be expensive and energy-intensive. CED reduces these costs by utilising lightweight hardware with efficient power consumption, enabling long-term deployment even in remote areas.
Our research aligns with global calls for leveraging technology to combat biodiversity loss and mitigate the impacts of human-induced environmental changes [smith2022]. It offers a significant step forward in integrating AI-driven solutions into ecological monitoring frameworks, bridging gaps between research, policy, and conservation practice. The rest of the paper is organised as follows. Section 2 reviews relevant literature. Section 3 outlines the methodology, including an overview of the CED prototype hardware, the datasets used, and the network architecture. Section 4 and 5 presents the training and experimental results, with a focus on model training results, stability and power consumption on the NJON. Finally, Section 6 concludes the paper and discusses directions for future work.
@@ 2 Related Work
Recent years have seen a significant increase in the interest in processing underwater images [mittal2022]. The interest is driven by the need to study and monitor aquatic plants and animals, which has applications in marine biology, economy, and biodiversity management. Understanding the behaviour and numbers of aquatic species is crucial for various purposes, including protecting endangered species, analysing differences in species, and early detection of climatic events like pollution and global warming. For instance, Plankton, which produces over 80% of the world’s oxygen [witman-2021], plays a critical role in the ocean’s food chain and atmosphere-ocean connection. Monitoring plankton levels is essential, as both low and excessive levels can have harmful consequences. Furthermore, organisms like Posidonia Oceanic, which thrive in clean water, are vital for biodiversity, beach erosion reduction, and water quality enhancement [mittal2022]. Image processing complements other techniques like physiochemical water analysis and sonar-based detection in understanding the impact of global warming and human activity on marine life.
@@ 2.1 The challenges of underwater monitoring
Underwater imaging contends with substantial energy loss, resulting in diminished and inconsistent illumination and visibility, particularly in the deeper expanses [mittal2022]. The situation is further exacerbated by the presence of freshwater and ocean currents, intensifying the degradation of visibility. Impurities, suspended solids, low contrast, and compromised edges and details further impact underwater images. The distortion of colour based on distance, caused by non-uniform spectral propagation, adds another layer of complexity. Additionally, underwater conditions present intricate and uncontrolled backgrounds, posing challenges in segregating subjects such as sea cucumbers and shellfish from their surroundings. The unbridled movement of fishes in three-dimensional space, often concealed behind objects, compounds the difficulty in accurately determining their pose, orientation, and size. The scarcity of large-scale, publicly available datasets for underwater images, coupled with imbalances, necessitates the incorporation of data augmentation techniques.
@@ 2.2 Early advances in underwater bio-diversity monitoring
In the early 90’s, [strachan-1990] laid the groundwork for automated fish grading based on size and colour, utilising image analysis to estimate fish shape and colour. The proposed approach allowed the differentiation of haddock fish stocks from the Rockall Plateau, located midway between Greenland and Iceland and the North Sea. The method involved digitised Red, Green, and Blue (RGB) images captured from beneath a conveyor belt, processed using bespoke software written in the C programming language. [strachan-1990] marked a significant advancement in distinguishing between fish species. Building on the pioneering work, [storbeck-2001] conducted further research in 2001 by integrating computer vision with a neural network, expanding on the foundations set by [strachan-1990]. Both studies required the fish to be laid flat on a conveyor belt moving at a specific speed, with a vertical arrangement of a laser source and camera above the belt. However, the classification method, while groundbreaking, proves impractical for underwater environments.
@@ 2.3 Modern application of computer vision for underwater bio-diversity monitoring
A more contemporary perspective is presented in the work of [Zhang], showcasing a modern application of computer vision for underwater object detection. Employing the YOLO v4 model, the authors achieved an mean Average Precision (mAP) of 92.65% at a processing speed of approximately 45 Frames per Second (FPS). Despite the success, it’s crucial to note that the achievement relied on a high-end NVIDIA graphics card for accelerating the YOLO v4. However, deploying such hardware alongside sensors in physical underwater settings becomes unfeasible due to size and power constraints. In response to the constraint, it is advisable to explore embedded systems as viable alternatives. Such platforms present a more pragmatic deployment option for sensors, ensuring the practical implementation of advanced computer vision techniques in underwater environments. As underwater computer vision emerges as a burgeoning field, studying analogous works becomes imperative to identify potential gaps and opportunities that can be leveraged to advance knowledge in this domain. For instance, [Hegde2021] successfully implemented a trained MobileNetV2 model capable of running on a Raspberry Pi, achieving an impressive precision of 98% for mackerel. While acknowledging that a YOLO model could offer faster results, they deemed it unfeasible due to hardware limitations. Consequently, their image capture and processing occurred at a rate of once every 2 seconds, equivalent to about 0.5 FPS. This limitation could prove critical in scenarios involving Underwater Autonomous Vehicles, where objects of interest might traverse the sensors before processing completion.
@@
The incorporation of a YOLO v5 model on the NJON platform is set to enhance performance, delivering a finely tuned solution tailored for the target applications. Noteworthy is the study by [Hegde2021], which had a limited implementation of comprehensive image pre-processing methods, mainly focusing on cropping. In contrast, our CED algorithm surpasses these limitations by integrating background subtraction which is a technique enables to extract moving objects (foreground) from the background, as discussed by [Huang2017] and [Kalsotra2021].
@@
The sophisticated enhancement is anticipated to streamline object identification, potentially fortifying our algorithm’s efficacy in detecting small objects [Hegde2021]. Importantly, our dataset boasts an extensive collection of 450 500 images for each category, significantly contributing to the robustness of our approach. In a comparable study, [Wang] employed YOLO nano on an NJON to accelerate the algorithm at the edge. The model used by [Wang] demonstrated the capability to detect underwater objects at an impressive frame rate of 8 FPS, accompanied by a notable mAP of approximately 74.8%. A discernible trade-off emerges between runtime and accuracy when juxtaposed with the work of [Hegde2021]. The strategic use of image pre-processing might provide an avenue for simpler, faster models to attain higher accuracy in underwater object detection scenarios.
@@
[Chen] presented a methodology similar to the one employed in this work, utilising data augmentation to enhance the capabilities of the object detection model. [Wei] proposed mosaic augmentation as use to augment the dataset to enhance the detection of smaller objects in cluttered environments. Moreover [Wei] introduced an innovative adaptation, involving the calculation of gains on the RGB channels to modify and adjust images, accommodating variations in underwater light levels. The CED distinguishes itself through the deliberate exclusion of the background.
@@
Considerable efforts have been directed towards advancing efficient autonomous underwater Underwater Autonomous Vehicles (UAVs) employing machine learning algorithms for debris detection. Leveraging the computational power of an NVIDIA Jetson TX2 GPU, [fulton-2019] conducted comprehensive experiments involving transfer learning on various models, including YOLO v2, Tiny-YOLO, Faster R-CNN, and Single Shot Detector (SSD). In their approach, the author fine-tuned each network individually for R-CNN, SSD, and Tiny-YOLO. However, for YOLO v2, an additional step involved training using transfer learning. The models were specifically tailored for underwater plastic debris detection, and the results indicated that Faster R-CNN achieved the highest accuracy, albeit with longer inference times. YOLO v2 struck a balance between speed and accuracy, SSD exhibited the best inference time, and Tiny-YOLO emerged as the top performer in terms of overall performance.
@@
[saleh-2024] asserts that automated fish phenotyping, encompassing parameters such as fish length, weight, size, and mortality, provides invaluable insights into understanding the comprehensive profile of classified fish. The task of individually counting fish is notably resource-intensive, demanding skilled labour for hours, thereby contributing to elevated project costs and extended delivery timelines. In a related domain, [ranjan-2023] introduced a real-time mortality monitoring system based on YOLO v7, achieving a commendable mAP of 93.4% and an F1 score of 0.89. The CED builds upon this foundation, not only incorporating these crucial findings, but also extending the phenotyping metrics. Specifically, we introduce age estimation, distinguishing between adult and juvenile crayfish.